How many ounces of water should a chemist remove from 20 ounces of a 20% sulfuric acid solution to increase it to a 40% sulfuric acid solution?
Answer by stanbon(75887)

How many ounces of water should a chemist remove from 20 ounces of a 20% sulfuric acid solution to increase it to a 40% sulfuric acid solution?
-------
Equation:
water - water = water
0.80*20 - x = 0.60(20-x)
-----
Multiply thru by 100 to get:
80*20 - 100x = 60*20 - 60x
40x = 20*20
x = 10 ounces (amt. of water to be removed)
----------------
Cheers,
Stan H.
========================